# FERMILAB-PUB-24-0295-T

### v1

@@ Abstract
The QCD axion and axion-like particles, as leading dark matter candidates, can also have interesting implications for dark matter substructures if the Peccei-Quinn symmetry is broken after inflation. In such a scenario, axion perturbations on small scales will lead to the formation of axion miniclusters at matter-radiation equality, and subsequently the formation of axion stars. Such compact objects open new windows for indirect searches for axions. We compute the axion star mass function based on recent axion minicluster studies and Bose star simulations. Applying this mass function, we find post-inflation axion-like particles with masses $m_{a}<3.3\times 10^{-17}$ eV are constrained by the lack of dynamical heating of stars in ultrafaint dwarfs. We also find that current microlensing surveys are insensitive to QCD axion stars. While we focus on the gravitational detectability of axion stars, our result can be directly applied to other interesting signatures of axion stars, e.g. their decay to photons, that require as input the abundance, mass, and density distribution of axion stars.
@@ 2 Overview of Axion Stars and Axion Minihalos
There are extensive studies on the stable solutions of axion stars (or generic Bose stars) [31, 32], the formation [33], and the collapse of axion stars [34, 35, 36]. However, the population analysis and detectable signatures are still lacking. Recently, there have been calculations to determine the axion star explosion rate [37, 29] and their encounter rates with neutron stars [38]. The mass function and density of axion minihalos, as the environments for axion star formation, are crucial to determining the formation rate and population of axion stars. In this work, we adopt the state-of-art axion minihalo population and axion star formation models to systematically study the axion star mass function. In what follows, we will give an overview of the current literature on axion star solutions, the evolution of axion minihalos, and the formation of axion stars. Applying these theoretical tools, we will derive a mass function of axion stars, which will be important inputs for indirect searches of axion dark matter.
@@ 2.1 Stable configurations of axion stars
In general, a lighter axion mass $m_{a}$ corresponds to a larger axion star due to the large de Broglie wavelength. On the other hand, a heavier axion star has a smaller radius, because it requires more kinetic pressure to balance the gravity. Overall, from Eq. 2.11 we see that the density of a star grows as the fourth power of its mass, $\rho_{\star}\sim M_{\star}^{4}(G_{N}m_{a}^{2})^{3}$. This will have important implications when we study the gravitational detectability of QCD axion stars. Since more compact objects can be probed with observations such as microlensing, one needs to form a sufficiently heavy axion star to reach a high density for detection. Therefore, the evolution and growth of axion stars will be a crucial part of the story (Unfortunately, as we will discuss later in Fig. 1, the QCD axion stars will not grow to a mass that will be compact enough for the current sensitivity of microlensing). As the mass of the axion star grows, Eq. 2.10 shows that at some point a sensible solution can no longer be obtained. This is due to the breakdown of the balance between gravity and gradient pressure caused by the non-negligible addition of attractive self-couplings, leading to the collapse of axion stars. The maximal mass that still corresponds to a stabilized axion star is
@@
The two solutions (2.10) merge at this critical mass (The $R_{\star}^{-}$ root is sometimes called the critical branch, which is metastable and not relevant to our study). After the collapse of critical mass stars, we will obtain the dense branch of axion stars which will soon emit relativistic axions [47, 48, 35] and are thus cosmologically unstable, with a lifetime $\sim 10^{3}m_{a}^{-1}$ [32]. In this work, we will not consider strong self-couplings that will destabilize axion stars. This corresponds to a large $f_{a}$ that is not excluded by Ref. [29, 30] that considers the collapse of axion stars and their observational constraints. In this work, we will focus on the stable axion stars and study their gravitational detectabilities.
@@ 2.2 An overview on axion minihalos
$\frac{df_{h}}{d{\rm ln}M_{h}}=\sqrt{\frac{\nu}{2\pi}}{\rm exp}(-\nu/2)\frac{d\,{\rm ln}\nu}{d\,{\rm ln}M_{h}}\,,$ (2.15)
where $M_{h}$ is the axion minicluster(or minihalo) mass and $df_{h}/d{\rm ln}M_{h}$ is the mass fraction of axion minihalos per logarithmic mass i.e. $f_{h}$ is the mass fraction of dark matter bound in minihalos. We choose $df_{h}/d{\rm ln}M_{h}$ to present the mass distribution of the population of axion minihalos, as it can characterize the most significant mass ranges of these objects that contain the majority of the total mass when considering logarithmic mass bins. $f_{h}=M_{h}n_{h}/\bar{\rho}_{a,0}$ is the mass density fraction of axion in minihalos. $\nu$ is defined as:
@@
As we see in the above expression, the only relevant parameter for $\sigma(M_{h})$ is the dimensionless quantity $x_{h}\equiv M_{h}/M_{h,c}$, which will give the full description of the dimensionless mass function in Eq. 2.15. At $\nu<1$, the exponential part of the mass function can be ignored and the mass function exhibits a power law $df_{h}/d{\rm ln}M_{h}\propto x_{h}^{1/2}$. As discussed before, we cut off the mass ranges $M_{h}<M_{h,c}$ (or $x_{h}<1$) to filter out the contributions from high wave numbers; we only consider $x_{h}\geq 1$.
@@ 2.3 Formation of axion stars in axion minihalos
Since the axion star configuration, as discussed earlier in this section, is an energetically favored state of axion fields, the axion waves prefer to reach the axion star configuration as long as there is enough time for relaxation. Such processes can be understood as Bose-Einstein condensation of Bosonic dark matter. Therefore, the relevant timescale of axion star formation is the scattering timescale between axions as wave-like particles, which approximately gives the condensation timescale of axion stars in simulations of axion fields [33, 85, 86, 87]
@@ 3 Axion Star Mass Functions
The first minihalo starts to form roughly at matter-radiation equality because the Hubble time at equality is approximately the same as the minihalo dynamical timescale needed to form a virialized minihalo. After the equality, axion stars can form and grow, but axion minihalos also grow and merge. As axion minihalos become heavier, the axion star formation time increases faster than the Hubble time, causing the growth of axion stars to cease. Also, we only consider one axion star per minihalo because they typically form in the center of minihalos, and if more than one stars are formed they quickly merge. Thus, we focus on the formation around matter-radiation equality, and the late evolution of axion stars such as mergers, which requires more extensive numerical studies, is not considered in this work. Our calculation underestimates the mass of individual stars, but we expect the total mass fraction in axion stars is well captured. This simple estimate already suggests that an $\mathcal{O}(1)$ fraction of axion dark matter forms axion stars, which can lead to interesting observational constraints and potential signals.
@@ 3.2 ALP Stars
$M_{h,c}$$\approx$ $14~{}M_{\odot}\left(\frac{10^{-16}~{}\mathrm{eV}}{m_{a}}\right)^{3/2}\left(\frac{g_{*}(T_{\rm osc})}{g_{*}(T_{0})}\right)^{1/4}~{},$ (3.9)
$M_{\star,c}$$\approx$ $37~{}M_{\odot}\,x_{h}^{1/12}\left(\frac{10^{-16}~{}\mathrm{eV}}{m_{a}}\right)^{3/2}\left(\frac{g_{*}(T_{\rm osc})}{g_{*}(T_{0})}\right)^{1/12}~{},$ (3.10)
$\tau H(z_{\rm eq})$$\approx$ $\frac{3.4\times 10^{-5}~{}x_{h}^{7/2}\left(\frac{g_{*}(T_{\rm osc})}{g_{*}(T_{0})}\right)^{1/2}}{\log\left[2.2x_{h}^{11/12}\left(\frac{g_{*}(T_{\rm osc})}{g_{*}(T_{0})}\right)^{1/6}\right]}~{}.$ (3.11)
Again, $T_{\rm osc}$ is determined by $m_{a}$ from Eq. 3.7. Here, we have ignored the small difference between $g_{*}(T)$ and $g_{*,S}(T)$. With the above expressions, we obtained all the equations needed to compute the axion star mass function for ALPs. Note that $\tau$ is independent of $m_{a}$ except for a minor dependence from $g_{*}$. This gives an almost universal mass distribution shape for all $m_{a}$, although as $m_{a}$ is increased the peak in the distribution of star masses also increases.
@@ 4 Constraints
Figure 5: Current status of constraints on ultralight dark matter. Here we compare our results with the existing constraints in the literature that apply to various dark matter models. For generic fuzzy dark matter, there are Lyman-$\alpha$ constraints ($m_{a}>2\times 10^{-21}$ eV [102]) and ultrafaint dwarf galaxies constraints ($m_{a}>3\times 10^{-19}$ eV [103]). Lyman-$\alpha$ also constrains post-inflationary axions ($m_{a}>3\times 10^{-18}$ eV [104]), which is the same scenario as discussed here, but the bound in this work is stronger by one-order-of-magnitude ($m_{a}>3.3\times 10^{-17}\rm eV$). There is also a generic bound on post-inflation dark matter ($m_{a}>10^{-19}$ eV) considering the axion isocurvature perturbations and the free-streaming effect, which will be again constrained by Lyman-$\alpha$ [105]. The constraint from ALP stars is derived by considering their dynamical heating effects in ultrafaint dwarfs [100, 101].
@@
where $f_{\textrm{UFD}}$ is the constraint on the mass fraction of compact objects assuming a single mass $M_{\star}$ and density $\rho_{\star}$ in Ref. [101]. Any axion star mass function that does not satisfy Eq. 4.1 will be ruled out. We can easily test if certain ALP masses are excluded by applying the axion star mass function presented in Fig. 4 to the above integral Eq. 4.1. Lighter axion masses are more constrained because there are stronger bounds for more massive objects. As a result, we exclude the ALP mass $m_{a}<3.3\times 10^{-17}~{}\mathrm{eV}$. We present our result in Fig. 5 in comparison to other constraints on fuzzy DM, post-inflation DM, and post-inflation axions. Our work focuses on the observational constraints of post-inflation axions, which feature an enhancement on the matter power spectrum as well as the formation of axion minihalos and axion stars. The constraint on axion minihalos only from Lyman-$\alpha$ has already been considered and leads to a weaker constraint [104]. There are also generic bounds on the mass of post-inflation dark matter [105] and fuzzy dark matter [102, 111, 103, 112], which are also weaker than our bound obtained from the gravitational effects of axion stars.
@@ Acknowledgement
We thank Peter Graham and Harikrishnan Ramani for the helpful discussions and for providing us with the data associated with their UFD constraint. We also thank Han Gil Choi, Wayne Hu, Xuheng Luo, and Dennis Maseizik for the useful discussions. JHC, PJF, and HX are supported by Fermi Research Alliance, LLC under Contract DE-AC02-07CH11359 with the U.S. Department of Energy.

### v3

@@ Abstract
The QCD axion and axion-like particles, as leading dark matter candidates, can also have interesting implications for dark matter substructures if the Peccei-Quinn symmetry is broken after inflation. In such a scenario, axion perturbations on small scales will lead to the formation of axion miniclusters at matter-radiation equality, and subsequently the formation of axion stars. Such compact objects open new windows for indirect searches for axions. We compute the axion star mass function based on recent axion minicluster studies and Bose star simulations. Applying this mass function, we find post-inflation axion-like particles with masses $1.8\times 10^{-21}~{}{\rm eV}<m_{a}<3.3\times 10^{-17}~{}{\rm eV}$ are constrained by the lack of dynamical heating of stars in ultrafaint dwarfs. We also find that current microlensing surveys are insensitive to QCD axion stars. While we focus on the gravitational detectability of axion stars, our result can be directly applied to other interesting signatures of axion stars, e.g. their decay to photons, that require as input the abundance, mass, and density distribution of axion stars.
@@ 2 Overview of Axion Stars and Axion Minihalos
There are extensive studies on the stable solutions of axion stars (or generic Bose stars) [31, 32, 33, 34, 35, 36, 37], the formation [38, 39, 40, 41], and the collapse of axion stars [42, 43, 44, 45]. However, the population analysis and detectable signatures are still lacking. Recently, there have been calculations to determine the axion star explosion rate [46, 29], the detection of axion star explosions [47, 48, 49, 50], and the encounter rate with axion minihalos and neutron stars [51]. The mass function and density of axion minihalos, as the environments for axion star formation, are crucial to determining the formation rate and population of axion stars. In this work, we adopt the state-of-art axion minihalo population and axion star formation models to systematically study the axion star mass function. In what follows, we will give an overview of the current literature on axion star solutions, the evolution of axion minihalos, and the formation of axion stars. Applying these theoretical tools, we will derive a mass function of axion stars, which will be important inputs for indirect searches of axion dark matter.
@@ 2.1 Stable configurations of axion stars
In general, a lighter axion mass $m_{a}$ corresponds to a larger axion star due to the large de Broglie wavelength. On the other hand, a heavier axion star has a smaller radius, because it requires more kinetic pressure to balance the gravity. Overall, from Eq. 2.11 we see that the density of a star grows as the fourth power of its mass, $\rho_{\star}\sim M_{\star}^{4}(G_{N}m_{a}^{2})^{3}$. This will have important implications when we study the gravitational detectability of QCD axion stars. Since more compact objects can be probed with observations such as microlensing, one needs to form a sufficiently heavy axion star to reach a high density for detection. Therefore, the evolution and growth of axion stars will be a crucial part of the story (Unfortunately, as we will discuss later in Fig. 1, the QCD axion stars will not grow to a mass that will be compact enough for the current sensitivity of microlensing). As the mass of the axion star grows, Eq. 2.10 shows that at some point a sensible solution can no longer be obtained. This is due to the breakdown of the balance between gravity and gradient pressure caused by the non-negligible addition of attractive self-couplings, leading to the collapse of axion stars. The maximal mass that still corresponds to a stabilized axion star is [56, 57, 60]
@@
The two solutions (2.10) merge at this critical mass (The $R_{\star}^{-}$ root is sometimes called the critical branch, which is metastable and not relevant to our study). After the collapse of critical mass stars, we will obtain the dense branch of axion stars which will soon emit relativistic axions [61, 62, 43] and are thus cosmologically unstable, with a lifetime $\sim 10^{3}m_{a}^{-1}$ [34, 45]. In this work, we will not consider strong self-couplings that will destabilize axion stars. This corresponds to a large $f_{a}$ that is not excluded by Ref. [29, 30] that considers the collapse of axion stars and their observational constraints. In this work, we will focus on the stable axion stars and study their gravitational detectabilities.
@@ 2.2 An overview on axion minihalos
$\frac{df_{h}}{d\log M_{h}}=\sqrt{\frac{\nu}{2\pi}}{\rm exp}(-\nu/2)\frac{d\log\nu}{d\log M_{h}}\,,$ (2.15)
where $M_{h}$ is the axion minicluster(or minihalo) mass and $df_{h}/d\log M_{h}$ is the mass fraction of axion minihalos per logarithmic mass i.e. $f_{h}$ is the mass fraction of dark matter bound in minihalos. We choose $df_{h}/d\log M_{h}$ to present the mass distribution of the population of axion minihalos, as it can characterize the most significant mass ranges of these objects that contain the majority of the total mass when considering logarithmic mass bins. $f_{h}=M_{h}n_{h}/\bar{\rho}_{a,0}$ is the mass density fraction of axion in minihalos. $\nu$ is defined as:
@@
As we see in the above expression, the only relevant parameter for $\sigma(M_{h})$ is the dimensionless quantity $x_{h}\equiv M_{h}/M_{h,c}$, which will give the full description of the dimensionless mass function in Eq. 2.15. At $\nu<1$, the exponential part of the mass function can be ignored and the mass function exhibits a power law $df_{h}/d\log M_{h}\propto x_{h}^{1/2}$. As discussed before, we cut off the mass ranges $M_{h}<M_{h,c}$ (or $x_{h}<1$) to filter out the contributions from high wave numbers; we only consider $x_{h}\geq 1$.
@@ 2.3 Formation of axion stars in axion minihalos
Since the axion star configuration, as discussed earlier in this section, is an energetically favored state of axion fields, the axion waves prefer to reach the axion star configuration as long as there is enough time for relaxation. Such processes can be understood as Bose-Einstein condensation of Bosonic dark matter. Therefore, the relevant timescale of axion star formation is the scattering timescale between axions as wave-like particles, which approximately gives the condensation timescale of axion stars in simulations of axion fields [38, 99, 100, 101, 102]
@@ 3 Axion Star Mass Functions
The first minihalo starts to form roughly at matter-radiation equality because the Hubble time at equality is approximately the same as the minihalo dynamical timescale needed to form a virialized minihalo [106, 107]. After the equality, axion stars can form and grow, but axion minihalos also grow and merge. As axion minihalos become heavier, the axion star formation time increases faster than the Hubble time, causing the growth of axion stars to cease. Also, we only consider one axion star per minihalo because they typically form in the center of minihalos, and if more than one stars are formed they quickly merge. Thus, we focus on the formation around matter-radiation equality, and the late evolution of axion stars such as mergers, which requires more extensive numerical studies, is not considered in this work. Our calculation underestimates the mass of individual stars, but we expect the total mass fraction in axion stars is well captured. This simple estimate already suggests that an $\mathcal{O}(1)$ fraction of axion dark matter forms axion stars, which can lead to interesting observational constraints and potential signals.
@@ 3.2 ALP Stars
$M_{h,c}$$\approx$ $3.1\,M_{\odot}\,\eta\left(\frac{10^{-16}~{}\mathrm{eV}}{m_{a}}\right)^{3/2}~{},$ (3.9)
$M_{\star,c}$$\approx$ $22\,M_{\odot}\,x_{h}^{1/12}\eta^{1/3}\left(\frac{10^{-16}~{}\mathrm{eV}}{m_{a}}\right)^{3/2}~{},$ (3.10)
$\tau H(z_{\rm eq})$$\approx$ $\frac{1.6\times 10^{-6}\,x_{h}^{7/2}\eta^{2}}{\log\left[0.78\,x_{h}^{11/12}\eta^{2/3}\right]}~{}.$ (3.11)
Here, $\eta\equiv g_{*S}(T_{\text{osc}})/g_{*}^{3/4}(T_{\text{osc}})$ and $T_{\rm osc}$ is determined by $m_{a}$ from Eq. 3.7. Note that the log factor in Eq. 3.11 is always positive. With the above expressions, we obtained all the equations needed to compute the axion star mass function for ALPs. Note that $\tau$ is independent of $m_{a}$ except for a minor dependence from $g_{*}$. This gives an almost universal mass distribution shape for all $m_{a}$, although as $m_{a}$ is increased the peak in the distribution of star masses also increases.
@@ 4 Constraints
Figure 5: Current status of constraints on ultralight dark matter. Here we compare our results with the existing constraints in the literature that apply to various dark matter models. For generic fuzzy dark matter, there are Lyman-$\alpha$ constraints ($m_{a}>2\times 10^{-21}$ eV [119]) and ultrafaint dwarf (UFD) galaxies constraints ($m_{a}>3\times 10^{-19}$ eV [120]). Lyman-$\alpha$ also constrains post-inflationary axions ($m_{a}>3\times 10^{-18}$ eV [121]), which is the same scenario as discussed here, but the bound in this work is stronger by one-order-of-magnitude ($m_{a}>3.3\times 10^{-17}\rm eV$). There is also a generic bound on post-inflation dark matter ($m_{a}>10^{-19}$ eV) considering the axion isocurvature perturbations and the free-streaming effect, which will be again constrained by Lyman-$\alpha$ [122]. The constraint from ALP stars is derived by considering their dynamical heating effects in ultrafaint dwarfs [117, 118].
@@
where $f_{\textrm{UFD}}$ is the constraint on the mass fraction of compact objects assuming a single mass $M_{\star}$ and density $\rho_{\star}$ in Ref. [118]. Any axion star mass function that does not satisfy Eq. 4.1 will be ruled out. We can easily test if certain ALP masses are excluded by applying the axion star mass function presented in Fig. 4 to the above integral Eq. 4.1. Lighter axion masses are more constrained because there are stronger bounds for more massive objects. As a result, we exclude the ALP mass $1.8\times 10^{-21}~{}{\rm eV}<m_{a}<3.3\times 10^{-17}~{}{\rm eV}$. We present our result in Fig. 5 in comparison to other constraints on fuzzy DM, post-inflation DM, and post-inflation axions. Our work focuses on the observational constraints of post-inflation axions, which feature an enhancement on the matter power spectrum as well as the formation of axion minihalos and axion stars. The constraint on axion minihalos only from Lyman-$\alpha$ has already been considered and leads to a weaker constraint [121]. There are also generic bounds on the mass of post-inflation dark matter [122] and fuzzy dark matter [119, 128, 120, 129], which are also weaker than our bound obtained from the gravitational effects of axion stars.
@@ Acknowledgement
We thank Peter Graham and Harikrishnan Ramani for the helpful discussions and for providing us with the data associated with their UFD constraint. We also thank Han Gil Choi, Wayne Hu, Xuheng Luo, and Dennis Maseizik for the useful discussions and Raymond Co for pointing out typos in some equations. JHC, PJF, and HX are supported by Fermi Research Alliance, LLC under Contract DE-AC02-07CH11359 with the U.S. Department of Energy.